[Hallc_running] Owl shift on 01/12 (tonight) on standby

Dipangkar Dutta ddutta07 at gmail.com
Thu Jan 11 14:03:26 EST 2018

Hi All,
The current plan at MCC is to setup 4-hall operation before anything else,
and the order in which the halls are being setup is D-A-B-C (because all
other halls are at 5 pass and we are at 3 pass). This means that the
chances of getting beam by the owl shift are very remote. The shift crew
during the swing shift please do the shift checklist at the beginning and
also if possible sometime close to the end of the shift, but you do not
have to stay for the whole shift. The owl shift tonight is being put on

The target cool-down also ran into some issues. Dave Meekins suspects a
vacuum problem in the return line, outside the hall. The cryo group is
investigating. We still plan on condensing on Friday during the day (has be
completed during regular hours).

We will lock up the Hall after the DOE tour is complete, around 4:00 PM and
go to power permit. The PD requests that we minimize the number of accesses
to the Hall after lock up.

I will send more updates if anything changes.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180111/35744c9d/attachment.html>

More information about the Hallc_running mailing list